OpenStudy (anonymous):

What is the solution of the system? 4x - 4y = -40 -4x + 7y = 55 (–5, –5) (5, 5) (5, –5) (–5, 5)

OpenStudy (anonymous):

What method for solving systems of equations do you know?

OpenStudy (anonymous):

i dont really know a method.. like i tried to turn 4x - 4y = -40 into 4x + 40/4 but it doesnt work

OpenStudy (anonymous):

Elimination would be a good way to do this one. So basically add together the two equations.

OpenStudy (anonymous):

@Acoustics95 first add both the equation u will get the value of y then put value of y in one of the equation given above u will get the required value of x

OpenStudy (anonymous):

how would i add the equations?

OpenStudy (anonymous):

Do it one piece at a time so 4x + -4x to start

OpenStudy (anonymous):

so x

OpenStudy (anonymous):

Almost, what is 4 - 4?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

so that would leave an x right?

OpenStudy (anonymous):

yes so you would have 0x, which is just 0.

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

now do the next pieces -4y + 7y

OpenStudy (anonymous):

3y

OpenStudy (anonymous):

ok, now what about the two items on the right sides of the =

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

good so now you have 3y=15, can you solve for y?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

good! now plug y=5 into one of the equations to solve for x

OpenStudy (anonymous):

4x - 20 = -40?

OpenStudy (anonymous):

good, keep going

OpenStudy (anonymous):

-4x + 30 = 55

OpenStudy (anonymous):

well you only have to use one of the equations to find x. so choose one to use your solving equation steps to find what x=

OpenStudy (anonymous):

so on the first equation the first 4x = -20?

OpenStudy (anonymous):

@rachelfreak123 ?

OpenStudy (anonymous):

yeah, now one more step to get the value

OpenStudy (anonymous):

whats the last step?

OpenStudy (anonymous):

divide both sides

OpenStudy (anonymous):

by -20?

OpenStudy (anonymous):

no, by the number in front of the x

OpenStudy (anonymous):

to the answer is (5, 5)?

OpenStudy (anonymous):

so*

OpenStudy (anonymous):

almost, you have 4x=-20, check your sign for x.

OpenStudy (anonymous):

oh so its 4/-20 so (5, -5)?

OpenStudy (anonymous):

well...... the x is negative, so remember that the values are in the order (x, y)

OpenStudy (anonymous):

im so confused.. :/

OpenStudy (anonymous):

well you know that the y=5 right? and the x=-5, so you just have to make srue you put it in the ( ) in the correct order

OpenStudy (phi):

4x = -20 4x/4 = -20/4 4/4 is 1 so that is x= -20/4 -20 divided by 4 is -5 x= -5

OpenStudy (anonymous):

ohhhhhhh so its (-5, 5)

OpenStudy (anonymous):

There you go!

OpenStudy (anonymous):

i see i mixed the places

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

Welcome